# Geometric calculations on probability manifolds from reciprocal relations in Master equations

### v3

@@ Abstract.
Onsager reciprocal relations model physical irreversible processes from complex systems. Recently, it has been shown that Onsager principles for master equations on finite states introduce a class of Riemannian metrics in a probability simplex, named probability manifolds. We call these manifolds finite-state generalized Wasserstein-$2$ spaces. In this paper, we study geometric calculations on probability manifolds, in which we derive the Levi-Civita connection, gradient, Hessian, and parallel transport, and compute the Riemannian and sectional curvatures. We present two examples of geometric quantities in probability manifolds. These include Levi-Civita connections from the chemical monomolecular triangle reaction and sectional, Ricci, and scalar curvatures in Wasserstein space on a three-point lattice.
@@ 1. Introductions
Non-equilibrium thermodynamics [4, 7] studies dynamical systems that interact with reservoirs for thermal energy or volume. Examples are coupled physical irreversible processes in complex systems, which are modeled by probability functions over physical states. Typical dynamics include thermoelectric phenomena and heat conduction in an anisotropic medium. To model irreversible processes, Onsager introduces the essential reciprocal relationship [21]. It describes the symmetric dissipation formulation of the probability transition equations arising from physical processes, known as free-energy dissipation.
@@
In recent years, a particular type of Onsager reciprocal relation has been widely studied in mathematics, known as gradient flows in optimal transport [2, 25, 17]. They characterize thermodynamically irreversible processes as gradient flows in probability space equipped with an infinite-dimensional Riemannian metric, namely Wasserstein-$2$ space. For example, gradient drift Fokker-Planck equations satisfy Onsager reciprocal relations, which are also gradient flows in Wasserstein-$2$ space [9, 22] for probability density functions on continuous states. Along this direction, master equations with detailed balance conditions [24] satisfy Onsager reciprocal relations. It has been shown that master equations can also be formulated as gradient flows in Wasserstein-$2$ spaces [5, 19, 20] for probability distributions on discrete states. From now on, we focus on generalized Wasserstein-$2$ spaces on a simplex set [21, 8], called thermodynamical probability manifolds or probability manifolds.
@@
Nowadays, geometric calculations in probability manifolds are essential for understanding and constructing fluctuation relations and thermodynamic properties. E.g., [15, 23] introduces overdamped Langevin dynamics in the probability manifold, namely Wasserstein common noises. It is based on geometric quantities, such as second-order operators in finite-state probability manifolds. A natural question arises. What are geometric quantities, such as Riemannian and sectional curvatures, in probability manifolds from Onsager reciprocal relations?
@@
In this paper, we study Riemannian calculations in a class of thermodynamical probability manifolds. We first compute the Levi-Civita connection, gradient, and Hessian operators on probability manifolds. We further derive Riemannian and sectional curvatures in the probability simplex set. Two examples of these computations are performed. One example is the Levi-Civita connection on a three-state probability simplex, where the discrete probability manifold is built from the master equation for a chemical monomolecular triangle reaction. The other example is the Ricci curvature on a probability manifold with a three-point lattice.
@@
In literature, geometric calculations in Wasserstein-$2$ spaces have been studied in [10, 18]. More recently, [11, 12, 13] generalized these calculations of metric spaces in both continuous and discrete states and studied divergence functions in Wasserstein-$2$ space. However, geometric calculations in generalized Wasserstein-$2$ metric spaces from the Onsager reciprocal relation on discrete states are still unknown. In this direction, [14, 16] has studied the Hessian operators in generalized metric spaces on continuous and discrete state spaces (lattices), respectively. In this paper, we formulate geometric calculations in thermodynamical probability manifolds based on generalized Onsager’s response matrices. We further compute Riemannian and sectional curvatures on probability manifolds. These curvatures are essential in understanding the high-order dissipation properties of free energies towards Wasserstein common noises [15, 23]. In the sequential work, we shall apply thermodynamical geometric quantities, such as the Ricci curvature lower bound, to study the convergence of stochastic Fokker-Planck equations on discrete states [15].
It is also worth mentioning that master equations and their Onsager reciprocal relations have been widely studied in the network theory [24]. Geometric calculations in probability manifolds also generalize this direction. The high-order derivative of Onsager response matrices defines the Riemannian curvature tensors in probability manifolds, shown in Theorem 2. We leave the network analysis of curvature tensors for future work.
@@
The paper is organized as follows. In section 2, we briefly review reciprocal relations, master equations with detailed balance conditions, and optimal transport metric spaces on discrete states. In section 3, we introduce the Riemannian geometric calculations, including Levi-Civita connections, parallel transport equations, and curvature tensors. We present an example of geometric calculations on a three-point probability simplex in section 4.
@@ 2. Onsager reciprocal relations, network notations, and Thermodynamical probability manifolds
In this section, we briefly review finite-state master equations with detailed balance relations, Onsager reciprocal relations, network notation, and a class of thermodynamical probability manifolds. See also detailed discussions in [19, 20, 24]. We briefly illustrate the equivalence relation between Wasserstein-type gradient flows and Onsager principles with generalized forces and response matrices; see [6].
@@ 2.1. Master equations with detailed balance relations
Consider a discrete state physical system $\big\{1,2,\cdots,n\big\}$. Denote a probability distribution $p(t)=(p_{i}(t))_{i=1}^{n}$ over states $i=1,2,\cdots,n$, that characterizes the discrete state system in a time domain $t\geq 0$, with
@@
where there is an initial value probability function $p(0)$, and the nonnegative quantity $Q_{ji}\geq 0$, $1\leq i\neq j\leq n$, is the constant transition probability per time from state $j$ to state $i$. The transition rate encodes the internal rate and external conditions, as determined by physical quantity differences between reservoir systems, such as temperature or pressure.
The master equation (1) describes a continuous time Markov process on a finite state $\big\{1,2,\cdots,n\big\}$ with a $Q$-matrix. Here $Q=(Q_{ij})_{1\leq i,j\leq n}$ forms the generator of the continuous time Markov process. Denote $\pi=(\pi_{i})_{i=1}^{n}\in\mathbb{R}^{n}$ with $\pi_{i}>0$ and $\sum_{i=1}^{n}\pi_{i}=1$ as a stationary point of equation (1). We also call $\pi$ as the invariant distribution satisfying
@@
A fact is that the master equation is with Onsager reciprocal relations, meaning that there exists a detailed balance condition for the Markov process with $Q$-matrix. This is also known as microscopic reversibility in thermal equilibrium; see an example in section 4. See also [24].
@@
Suppose that there exists a vector $\pi=(\pi_{i})_{i=1}^{n}\in\mathbb{R}^{n}$, with $\pi_{i}>0$ and $\sum_{i=1}^{n}\pi_{i}=1$, such that $Q$ and $\pi$ satisfy the following condition:
@@ 2.2. Onsager reciprocal relations
From Onsager reciprocal relations, one can rewrite equation (1) into a symmetric dissipation formulation. We shortly illustrate that this formulation represents gradient flows of free energies on probability manifolds.
@@
From the detailed balance condition (2), the weighted matrix $\omega$ is symmetric, since $\omega_{ij}=\omega_{ji}=Q_{ji}\pi_{j}=Q_{ij}\pi_{i}$. For a convex function $f\in C^{1}(\mathbb{R};\mathbb{R})$ with $f(1)=0$, master equation (1) can be formulated as below:
$\begin{split}\frac{dp_{i}(t)}{dt}=&\sum_{j=1}^{n}\Big[Q_{ji}\pi_{j}\frac{p_{j}(t)}{\pi_{j}}-Q_{ij}\pi_{i}\frac{p_{i}(t)}{\pi_{i}}\Big]\\ =&\sum_{j=1}^{n}\omega_{ij}\Big(\frac{p_{j}(t)}{\pi_{j}}-\frac{p_{i}(t)}{\pi_{i}}\Big)\\ =&\sum_{j=1}^{n}\omega_{ij}\frac{\frac{p_{j}(t)}{\pi_{j}}-\frac{p_{i}(t)}{\pi_{i}}}{f^{\prime}(\frac{p_{j}(t)}{\pi_{j}})-f^{\prime}(\frac{p_{i}(t)}{\pi_{i}})}\cdot\big(f^{\prime}(\frac{p_{j}(t)}{\pi_{j}})-f^{\prime}(\frac{p_{i}(t)}{\pi_{i}})\big)\\ =&\sum_{j=1}^{n}\omega_{ij}\theta_{ij}(p(t))\big(f^{\prime}(\frac{p_{j}(t)}{\pi_{j}})-f^{\prime}(\frac{p_{i}(t)}{\pi_{i}})\big).\end{split}$ (4)
@@
In fact, the R.H.S. of symmetric formulation (4) can be written as a matrix function times a vector function. This is known as the strong Onsager gradient flow. In physics, they represent a symmetric matrix function (Onsager’s response matrix) and the gradient of a free-energy function (the force). To illustrate this point, we first define a force vector function.
@@
For simplicity of notation, we also denote $L(\theta)=L(\theta(p))$. We note that the matrix $L(\theta)$ is symmetric and semi-positive definite. I.e.,
@@
where $\nabla_{p}\mathrm{D}_{f}(p\|\pi)$ is the generalized force and $L(\theta)$ is the Onsager’s response matrix. In addition, along with the solution of the master equation (1), the free energy $\mathrm{D}_{f}(p(t)\|\pi)$ decays in terms of the time variable.
@@
Example 1 (KL divergence mean).
@@
We note that Onsager reciprocal relations [21] are not limited to equation (1). These relations show that the macroscopic physical evolution equation can be written in the form (6) for general choices of weight functions $\theta$ and $f$-divergences (free energies). E.g., one can derive a general weight function $\theta$ from nonlinear master equations. On the other hand, many scientific computing methods also derive new classes of weight functions. Typical examples include the geometric mean, Harmonic mean, and arithmetic mean. See [6, 19, 20]. These mean functions are also studied in modeling dynamical density functional theory [26].
@@
We remark that the geometric calculations hold for general choices of average functions $\theta$.
@@ 2.3. Network notations
Formulation (8) is known as the Wasserstein gradient flow on finite state spaces [2, 19, 20]. In addition, the dissipation of free energy in (9) equals the squared norm in the probability simplex with optimal-transport-induced Riemannian metrics on finite states. We present these properties in the next section.
@@ 2.4. Thermodynamical probability manifolds
We remark that the map $\Phi\rightarrow\mathsf{V}_{\Phi}$ is an isomorphism $\mathbb{R}^{n}/\mathbb{R}\rightarrow T_{p}\mathcal{P}_{+}$. This is true from the diagonalization of $L(\theta)$ in (10), since there is only one constant eigenvector $u_{0}(p)$ with eigenvalue 0 for any probability $p\in\mathcal{P}_{+}$.
@@
We next illustrate some basic quantities in the manifold $(\mathcal{P}_{+},\mathbf{g})$. We first study the gradient operator. Mathematically, the Onsager reciprocal relation shows that the master equation (1) is a gradient descent flow of ff–divergence in $(\mathcal{P}_{+},\mathbf{g})$.
Denote $\bar{\mathrm{grad}}=\mathrm{grad}_{\mathbf{g}}\colon\mathcal{P}_{+}\times C^{\infty}(\mathcal{P}_{+};\mathbb{R})\rightarrow T_{p}\mathcal{P}_{+}$.
@@
When $F(p)=\mathrm{D}_{f}(p\|\pi)$, from the definition of $\bar{\mathrm{grad}}F$ and the R.H.S. of (4), we prove the equation (11). ∎
@@
where the minimization is taken over all continuous curves in the probability simplex set, $\gamma(t)\in C([0,1];\mathcal{P}_{+})$, $t\in[0,1]$, connecting boundary points $\gamma(0)=p^{0}$, $\gamma(1)=p^{1}$. We note that the minimal value function $\mathrm{Dist}=\mathrm{Dist}_{\mathbf{g}}\colon\mathcal{P}_{+}\times\mathcal{P}_{+}\rightarrow\mathbb{R}$ represents the Wasserstein-$2$ distance on the simplex set.

### v4

@@ Abstract.
Onsager reciprocal relations model physical irreversible processes from complex systems. Recently, it has been shown that Onsager principles for master equations on finite states introduce a class of Riemannian metrics on the probability simplex, leading to probability manifolds or finite-state Wasserstein–2 spaces. In this paper, we study geometric calculations on probability manifolds, deriving the Levi-Civita connection, gradient, Hessian operators of energies, parallel transport, and calculating both the Riemannian and sectional curvatures. We present two examples of geometric quantities in probability manifolds. One example is the Levi-Civita connection from the chemical monomolecular triangle reaction. The other example is the sectional, Ricci, and scalar curvatures in Wasserstein space on a three-point lattice graph.
@@ 1. Introductions
Non-equilibrium thermodynamics [5, 8] studies dynamical systems that interact with reservoirs for thermal energy or volume. Examples are coupled physical irreversible processes in complex systems, which are modeled by probability functions over physical states. Typical dynamics include thermoelectric phenomena and heat conduction in an anisotropic medium. In the modeling of irreversible processes, Onsager introduces the reciprocal relationship [22], which describes the symmetric dissipation of the probability transition equations arising from physical processes, known as free-energy dissipation or entropy production. Here, the free energy dissipation follows the second law of thermodynamics. Rather than addressing arbitrary physical mechanisms, one can focus on two canonical classes of stochastic models: overdamped Langevin dynamics in continuous state spaces, and detailed-balance Markov processes with transition rates on discrete domains. In both cases, entropy production is expressed as the time derivative of an entropy (or free-energy) functional along the governing probability evolution equations, such as the Fokker–Planck equation in the continuous setting or the master equation in the discrete setting.
@@
In recent decades, a particular type of gradient flow has been widely studied in optimal transport theory [2, 26, 18]. They characterize a class of partial differential equations as gradient flows of energies on a probability space equipped with an infinite-dimensional Riemannian metric, namely the Wasserstein-$2$ space. Among these dynamics, the famous example is the probability density transition equation for overdamped Langevin dynamics, which is the gradient-drift Fokker-Planck equation. In this setting, the gradient drift Fokker-Planck equation satisfies the gradient flow of free energy in Wasserstein-$2$ space [10, 23]. From the angle of non-equilibrium thermodynamics [9], the gradient drift Fokker-Planck equation satisfies the Onsager reciprocal relation. Along this direction, master equations with detailed balance conditions [25] can also be formulated as gradient flows in discrete Wasserstein-$2$ spaces [6, 20, 21] for probability distributions on discrete states. It also satisfies Onsager reciprocal relations [22]. From now on, we mainly study these generalized Wasserstein-$2$ spaces on a simplex, called thermodynamical probability manifolds, in short probability manifolds.
@@
Nowadays, geometric calculations in probability manifolds are essential for understanding fluctuation relations and thermodynamic properties. E.g., [16, 24] introduces overdamped Langevin dynamics in the probability manifold, namely Wasserstein common noises. It is based on geometric quantities, such as second-order operators in finite-state probability manifolds. Along this direction, a natural question arises. What are geometric quantities, including Riemannian and sectional curvatures, in probability manifolds?
@@
In this paper, we study Riemannian calculations in a class of thermodynamical probability manifolds. We first compute the Levi-Civita connection, gradient, and Hessian operators on probability manifolds. We further derive Riemannian and sectional curvatures in the probability simplex set. Two examples of these computations are performed. One example is the Levi-Civita connection on a three-state probability simplex, where the discrete probability manifold is constructed from the master equation for a chemical monomolecular triangle reaction. The other example is the Ricci curvature on a probability manifold with a three-point lattice graph.
@@
In literature, geometric calculations in Wasserstein-$2$ spaces have been studied in [11, 19]. More recently, [12, 13, 14] generalized these calculations of metric spaces in both continuous and discrete states and studied divergence functions in Wasserstein-$2$ space. However, geometric calculations in generalized Wasserstein-$2$ metric spaces from the Onsager reciprocal relation on discrete states are still unknown. In this direction, [15, 17] has studied the Hessian operators in generalized metric spaces on continuous and discrete state spaces (lattices), respectively. In this paper, we formulate geometric calculations in thermodynamical probability manifolds based on generalized Onsager’s response matrices. We further compute Riemannian and sectional curvatures on probability manifolds. It is also worth mentioning that master equations and their Onsager reciprocal relations have been widely studied in the network theory [25]. Geometric calculations in probability manifolds also generalize this direction. The high-order derivative of Onsager response matrices defines the Riemannian curvature tensors in probability manifolds, shown in Theorem 2. We leave the related network analysis of curvature tensors for future work.
In addition, there exist several studies connecting Onsager reciprocal relations [9], information geometry [28], optimal transport on graphs [6, 20, 21], and dynamical density functional theory [27]. In particular, [28] investigates a class of Onsager response matrices constructed from the Hessian of entropy functionals and interprets Onsager gradient flows as natural gradient flows in information geometry. In the present work, the Onsager response matrix is given from the mean function $\theta$ arising in master equations and depends on the choice of free energy. From this perspective, intrinsic geometric properties of probability manifolds–such as connections and curvatures–can be related to Hessian and higher-order derivatives of the free energy, objects that are central in information geometry [1, 3, 29]. See a concrete example in Example 5.
@@
The paper is organized as follows. In section 2, we briefly review reciprocal relations, master equations with detailed balance conditions, and optimal transport metric spaces on discrete states. In section 3, we introduce the Riemannian geometric calculations, including Levi-Civita connections, parallel transport equations, and curvature tensors. We also present two examples of geometric calculations on a three-point probability simplex in sections 4 and 5.
@@ 2. Onsager reciprocal relations, network notations, and Thermodynamical probability manifolds
This section reviews finite-state master equations with detailed balance relations, Onsager reciprocal relations, network notation, and a class of thermodynamical probability manifolds. These studies review results in [20, 21, 25].
@@ 2.1. Master equations with detailed balance relations
Consider a discrete state $\big\{1,2,\cdots,n\big\}$. Denote a probability distribution $p(t)=(p_{i}(t))_{i=1}^{n}\in\mathbb{R}^{n}_{+}$ over states $i=1,2,\cdots,n$, that characterizes the discrete state system in a time domain $t\geq 0$, with
@@
where there is an initial value probability function $p(0)$, and the nonnegative quantity $Q_{ji}\geq 0$, $1\leq i\neq j\leq n$, is the constant transition probability per time from state $j$ to state $i$. In modeling, this transition rate often represents the internal rate and external conditions, as determined by physical quantity differences between reservoir systems, such as temperature or pressure.
Master equation (1) describes a continuous time Markov process on a finite state $\big\{1,2,\cdots,n\big\}$ with a $Q$-matrix. Here, matrix $Q=(Q_{ij})_{1\leq i,j\leq n}$ is the generator of the continuous time Markov process. Denote $\pi=(\pi_{i})_{i=1}^{n}\in\mathbb{R}^{n}$ with $\pi_{i}>0$ and $\sum_{i=1}^{n}\pi_{i}=1$ as a stationary point of equation (1). We also call $\pi$ as the invariant distribution satisfying
@@
A fact is that the master equation is with Onsager reciprocal relations, meaning that there exists a detailed balance condition for the Markov process with $Q$-matrix. See an example in section 4. More examples have been studied in [25].
@@
Suppose that there exists a vector $\pi=(\pi_{i})_{i=1}^{n}\in\mathbb{R}^{n}$, with $\pi_{i}>0$ and $\sum_{i=1}^{n}\pi_{i}=1$, such that
@@ 2.2. Onsager reciprocal relations
One can rewrite equation (1) into a symmetric dissipation format [6, 20, 21], which satisfy Onsager reciprocal relations. This formulation represents gradient flows of free energies on probability manifolds.
@@
From the detailed balance condition (2), the weighted matrix $\omega$ is symmetric, since $\omega_{ij}=Q_{ij}\pi_{i}=Q_{ji}\pi_{j}=\omega_{ji}$. For a convex function $f\in C^{1}(\mathbb{R};\mathbb{R})$ with $f(1)=0$, master equation (1) can be written as below:
$\begin{split}\frac{dp_{i}(t)}{dt}=&\sum_{j=1}^{n}\Big[Q_{ji}\pi_{j}\frac{p_{j}(t)}{\pi_{j}}-Q_{ij}\pi_{i}\frac{p_{i}(t)}{\pi_{i}}\Big]=\sum_{j=1}^{n}\omega_{ij}\Big(\frac{p_{j}(t)}{\pi_{j}}-\frac{p_{i}(t)}{\pi_{i}}\Big)\\ =&\sum_{j=1}^{n}\omega_{ij}\frac{\frac{p_{j}(t)}{\pi_{j}}-\frac{p_{i}(t)}{\pi_{i}}}{f^{\prime}(\frac{p_{j}(t)}{\pi_{j}})-f^{\prime}(\frac{p_{i}(t)}{\pi_{i}})}\cdot\big(f^{\prime}(\frac{p_{j}(t)}{\pi_{j}})-f^{\prime}(\frac{p_{i}(t)}{\pi_{i}})\big)\\ =&\sum_{j=1}^{n}\omega_{ij}\theta_{ij}(p(t))\big(f^{\prime}(\frac{p_{j}(t)}{\pi_{j}})-f^{\prime}(\frac{p_{i}(t)}{\pi_{i}})\big).\end{split}$ (4)
@@
In fact, the R.H.S. of symmetric formulation (4) can be written as a matrix function times a vector function. This is known as the strong Onsager gradient flow (physical terminology), which contains a symmetric matrix function (Onsager’s response matrix) and the gradient of a free-energy function (the force). To illustrate this point, we first define a force vector function.
@@
For simplicity of notation, we write $L(\theta)=L(\theta(p))$. We note that the matrix $L(\theta)$ is symmetric and semi-positive definite. I.e.,
@@
where $\nabla_{p}\mathrm{D}_{f}(p\|\pi)$ is the generalized force and $L(\theta)$ is the Onsager’s response matrix. In addition, along with the solution of the master equation (1), the free energy $\mathrm{D}_{f}(p(t)\|\pi)$ decays in the time variable.
@@
Example 1 (KL divergence mean [20, 21]).
@@
Remark 1.
We note that Onsager reciprocal relations [22] are not limited to equation (1). These relations show that the macroscopic physical evolution equation can be written in the form (6) for general choices of weight functions $\theta$ and $f$-divergences (free energies). E.g., one can derive a general weight function $\theta$ from nonlinear master equations. On the other hand, many scientific computing methods also derive new classes of weight functions. Typical examples include the geometric mean, Harmonic mean, and arithmetic mean. See [7, 20, 21]. These mean functions are also studied in modeling dynamical density functional theory [27]. The geometric calculations in this paper hold for general choices of average functions $\theta=(\theta_{ij})_{(i,j)\in E}$, under the assumption that $\theta_{ij}\in C^{2}(\mathcal{P}_{+})$, for any indices $(i,j)\in E$.
@@ 2.3. Network notations
Equation (8) is known as the Wasserstein gradient flow on finite state spaces [2, 20, 21]. In addition, the dissipation of free energy in (9) equals the squared norm in the probability simplex with optimal-transport-induced Riemannian metrics on finite states. We present these properties in the next section.
@@ 2.4. Thermodynamical probability manifolds
We remark that the map $\Phi\rightarrow\mathsf{V}_{\Phi}$ is an isomorphism $\mathbb{R}^{n}/\mathbb{R}\rightarrow T_{p}\mathcal{P}_{+}$. This is true from the diagonalization of $L(\theta)$ in (10), since there is only one constant eigenvector $u_{0}(p)$ with eigenvalue 0 for any vector $p\in\mathcal{P}_{+}$.
@@
We next illustrate some basic quantities in the manifold $(\mathcal{P}_{+},\mathbf{g})$. We first study the gradient operator. Mathematically, the Onsager reciprocal relation shows that master equation (1) is a gradient descent flow of ff–divergence in $(\mathcal{P}_{+},\mathbf{g})$. Denote $\bar{\mathrm{grad}}=\mathrm{grad}_{\mathbf{g}}\colon\mathcal{P}_{+}\times C^{\infty}(\mathcal{P}_{+};\mathbb{R})\rightarrow T_{p}\mathcal{P}_{+}$.
@@
When $F(p)=\mathrm{D}_{f}(p\|\pi)$, from the definition of $\bar{\mathrm{grad}}F$ and the R.H.S. of (4), we derive equation (11). ∎
@@
where the minimization is taken over all continuous differentiable curves in the probability simplex set, $\gamma(t)\in C^{1}([0,1];\mathcal{P}_{+})$, $t\in[0,1]$, connecting boundary points $\gamma(0)=p^{0}$, $\gamma(1)=p^{1}$. We note that the minimal value function $\mathrm{Dist}=\mathrm{Dist}_{\mathbf{g}}\colon\mathcal{P}_{+}\times\mathcal{P}_{+}\rightarrow\mathbb{R}$ represents the Wasserstein-$2$ distance on the simplex set.